Calculate the voltage at points A, B, and C, in the following circuit. Find the power dissipated by each resistor. 6 V Calculate the total current I flowing through the following circuit segment. Then, find the voltage drop across each resistor. -18 V 12 kΩ 12 k
Answer #1

(a)

В 0 -V We haw awst amen 1 к.ru.vnyme,

Power dissipated across 10 kOmega resistor,

P10 = I2R = (0.6*10-3)2*10 * 103 = 0.36 * 10-2 W

      = 3.6 *10-3 W

Power dissipated across 1 kOmega resistor,

P1 = I2R = (3*10-3)2*1 * 103 = 9 * 10-3 W

      = 9 *10-3 W

(b)

금 + 3 2 12 12 4노고 2. AR R24 2-0

Now,

From figure 3 of the circuit,

Current across 4Omega resistor , i1 = (5 / (5 + 4) ) * 5.4

                                                 = 5/9 * 5.4 = 3 mA

Current across 5Omega resistor , i2 = 2.4 mA

Now,

figure 2 of the circuit we can see, Current is same in 4kOmega and 1 kOmega resistors as they are in series. i.e 2.4 mA

Now,

In figure 1 of the circuit we can see, 4kOmega resistor is made of three 12 kOmega resistors connected in parallel.

Thus, current will be equally divided into the 12 kOmega resistors.

=> i3 = 2.4 / 3 = 0.8 mA

Thus,

Current across all the 12 kOmega resistors is 0.8 mA

Current across 1 kOmega resistors is 2.4 mA

Current across the 4 kOmega resistors is 3 mA